Jul 31, 2014 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11:00 a.m. Eastern

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Tentative Agenda

  1. Announcements
  2. Action Items
  3. FCAL HDFCAL log book
  4. BCAL HDBCAL log book
  5. Reconstruction and Simulation
    1. Simulation of Spectra for Cosmic Muons in BCAL (Andrei & Irina)
  6. Calibration
  7. Any other business


Attendees: Mark D., Will M., Laura, Shayne, Manuel (JLab), Matt (IU), Sean (NWU), Tegan, Andrei, Zisis (Regina)

  1. Announcements: Magnet is cooling down nicely. It will run at 1200 A for about a week and then it will be deliberately dumped.
  2. Action Items: too many to list :-)
  3. FCAL: Manuel showed recent (today) logbook entries for base monitoring and pedestals. The former shows strange behavior for a cluster of four PMTs. Several tests were done and things look ok but Manuel will keep an eye on these. The pedestal plots in red show a problem. The pedestals appear at channel 2000 instead of 100. Mark mentioned that something similar was seen in the BCAL and the solution is to power-cycle the crates. Manuel will check this.
  4. BCAL: Zisis reported on Tuesday's meeting between JLab and Regina, taking advantage of Elton's presence. All students are proceeding with their projects. Laura and Shayne are carrying out more cosmics runs as well as building two separate two Monte Carlos to study the cosmics rates and energy depositions, respectively. Laura provided a table with the cosmics runs, that shows workable stats for the North side but the South has poor statistics. Once we get control back from Alex due to DAQ work, we have to address this deficit. Will is looking at FADC modes e.g. pulse integral, timing etc) and working on scripts to extract the Michel electron information. The next big tasks is to "commission" the TDCs. Mark reports that the Electronics group has upgraded the relevant systems to accommodate different trigger types, configurations, and updated procedures, so that things are easier to use now. Tests will commence after lunch. Will is modelling the TDC scripts after the ADC ones, so that we can see channels "in the green", problem ones etc.
  5. Reconstruction and simulation: Andrei presented preliminary results from the MC simulation on the fine geometry (energy depositions in the fiber volumes). Shoulders are visible on slide 3, due to the fiber structure. He stressed that the means deviate considerably from expectations on contrast to peak positions (from fitting) that nicely track the energy depositions based on the thicknesses of the cells for Module 13 (12 o'clock) and Module 1 (9 o'clock). When Irina returns to Regina, the 3.75 deg rotation, the 14" top trigger offset will be implemented and all simulations rerun. Expect this in 2-3 weeks and numbers will be good then, albeit without magnet yoke in the simulations, etc. Next step would be to take cosmics using the bias distribution system, so as to isolate individual SiPMs within a sum. Zisis asked for the widths of the peaks to be extracted in the tables for ccdb referencing in the future. Andrei asked for ROOT files with the module coordinate indexing. Mark will send ROOT files with both crate/slot/channel indexing and module indexing, stored in two separate trees. This will allow Andrei to extract the trigger scintillator ADC info from the new crate through the crate tree and apply it to the module tree, referenced by event no key.
  6. Calibration: Sean reported that the software update for lower level corrections to the data in accordance with what the electronics give is ongoing. The calorimeters look fine in this exercise but other detectors have issues. Sean also requested the Calorimetry Working Group people attend the Calibration meetings, and Matt and Zisis agreed to do so.
  7. AOB: none.